OpenStudy (anonymous):

Jim Moore opened a new savings account. He deposited $12,000 at 12 percent compounded semiannually. At the start of the fourth year, Jim deposits an additional $50,000 that is also compounded semiannually at 12 percent. At the end of 6 years, the balance in Jim Moore's account is: (Use the tables in the handbook)

OpenStudy (anonymous):

\[12000\times ( 1+.06)^{12} +50000\times ( 1+.06)^{8} \]$103,838.76

OpenStudy (anonymous):

It is 12000 x (1.06)^12 + 50000 x (1.06)^6 = 95,072.31 start of 4th year to end of 6th year = 6 semi-annual periods where interest is compounded for the second deposit